# Bogomolov property and Galois representations

### v1

@@ 1. Introduction
    Let $j$ and $i$ be integers with $1\leq j\leq n$ and $q^{j-1}\leq i\leq q^{j}-1$.The higher ramification groups are given by
@@
    In particular, the last non trivial higher ramification group has index $q^{n-1}-1$ and
    $G_{q^{n-1}-1}=\begin{cases}({\mathbb{Z}}/p{\mathbb{Z}})^{2}&\hbox{if }n\geq 2\\ {\mathbb{Z}}/(q-1){\mathbb{Z}}&\hbox{if }n=1.\end{cases}$
@@
    If $M$ is an integer prime to $p$ and $v\mid p$, then the image of ${\rho_{v}}_{|G_{{\mathbb{Q}}_{q}}}$ contains the scalar matrix $\left(\begin{smallmatrix}M&0\\ 0&M\end{smallmatrix}\right)$.
@@
$\rho_{v|G_{{\mathbb{Q}}_{q}}}\simeq\left(\mathrm{ind}_{G_{{\mathbb{Q}}_{q}}}^{G_{{\mathbb{Q}}_{p}}}\epsilon_{2}\right)\otimes\mu_{\sqrt{-1}}$
where $\epsilon_{2}:G_{{\mathbb{Q}}_{q}}\to{\mathbb{Z}}_{q}^{\times}$ is the cyclotomic characters and $\mu_{\sqrt{-1}}$ is a finite character of $G_{{\mathbb{Q}}_{q}}$ which take as values the fourth roots of unity. By Mackey’s restriction formula, we see that
(4.4) $\left(\mathrm{ind}_{G_{{\mathbb{Q}}_{q}}}^{G_{{\mathbb{Q}}_{p}}}\epsilon_{2}\right)_{|G_{{\mathbb{Q}}_{q}}}=\epsilon_{2}\oplus\epsilon^{\prime}_{2},$
where $\epsilon^{\prime}_{2}$ is the character obtained from $\epsilon_{2}$ by composing with the algebraic conjugation of ${\mathbb{Q}}_{q}/{\mathbb{Q}}_{p}$. Let for short $\omega=\mu_{\sqrt{-1}|G_{{\mathbb{Q}}_{q}}}$ and remark that $\omega$ takes values in {±$\{\pm 1\}$ and thus it is stable by the algebraic conjugation of ${\mathbb{Q}}_{q}/{\mathbb{Q}}_{p}$. Then
(4.5) $\rho_{v|G_{{\mathbb{Q}}_{q}}}\hbox{ is conjugate to }\qquad\begin{aligned} G_{{\mathbb{Q}}_{q}}&\longrightarrow\mathrm{GL}_{2}({\mathbb{Z}}_{q})\\ \sigma&\to\left(\begin{smallmatrix}(\epsilon_{2}\omega)(\sigma)&0\\ 0&(\epsilon_{2}\omega)^{\prime}(\sigma)\end{smallmatrix}\right).\end{aligned}$
@@
$\ker\rho(p^{n})_{|G_{{\mathbb{Q}}_{q}}}$ =$=\{\sigma\in G_{{\mathbb{Q}}_{q}}\ |\ (\epsilon_{2}\omega)(\sigma)\equiv(\epsilon_{2}\omega)^{\prime}(\sigma)\equiv 1\pmod{p^{n}}\}$
=$=\{\sigma\in G_{{\mathbb{Q}}_{q}}\ |\ (\epsilon_{2}\omega)(\sigma)\equiv 1\pmod{p^{n}}\}=\ker(\nu_{n}\omega_{n}),$
where we have denoted by $\omega_{n}$, $\nu_{n}$ the composition of $\omega$, $\nu$ respecively with the reduction modulo $p^{n}$. Thus
@@
By [9, Théorème I.2.1], $\nu$ is up to isogeny the Tate module of a Lubin–Tate formal group $\mathcal{F}$ over ${\mathbb{Z}}_{q}$. Then ${\mathbb{Q}}_{q}(\nu_{n})={\mathbb{Q}}_{q}(\mathcal{F}(n))$, where $\mathcal{F}(n)$ is the group of $p^{n}$-division points of $\mathcal{F}$. It follows from [19, Theorem V.5.4] and [19, Proposition V.6.1], that
  1. $\bullet$
    The extension ${\mathbb{Q}}_{q}(\nu_{n})/{\mathbb{Q}}_{q}$ is totally ramified and abelian of degree $(q-1)q^{n-1}$, and
    $\mathrm{Gal}({\mathbb{Q}}_{q}(\nu_{n})/{\mathbb{Q}}_{q})\simeq\left({\mathbb{Z}}_{q}/p^{n}{\mathbb{Z}}_{q}\right)^{\times}\simeq{\mathbb{Z}}/(q-1){\mathbb{Z}}\times\left({\mathbb{Z}}/p^{n-1}{\mathbb{Z}}\right)^{2}.$
  2. $\bullet$
    For integers $i,j$ such that $1\leq j\leq n$ and $q^{j-1}\leq i\leq q^{j}-1$, the higher ramification groups of ${\mathbb{Q}}_{q}(\nu_{n})$ are given by
    $G_{i}({\mathbb{Q}}_{q}(\nu_{n})/{\mathbb{Q}}_{q})=\mathrm{Gal}({\mathbb{Q}}_{q}(\nu_{n})/{\mathbb{Q}}_{q}(\nu_{j}))$
By its definition, $\nu$ is totally ramified and surjective; on the other hand $\omega$ is finite. By (4.6) and by Lemma 4.5 (iii) with $F={\mathbb{Q}}_{q}$ and with $\nu$, $\omega$ replaced by $\nu_{n}$, $\omega_{n}$ respectively
$\mathrm{Gal}({\mathbb{Q}}_{q}(p^{n})/{\mathbb{Q}}_{q})=\mathrm{Gal}({\mathbb{Q}}_{q}(\nu_{n}\omega_{n})/{\mathbb{Q}}_{q})\simeq\mathrm{Gal}({\mathbb{Q}}_{q}(\nu_{n})/{\mathbb{Q}}_{q})$
and
$G_{i}({\mathbb{Q}}_{q}(p^{n})/{\mathbb{Q}}_{q})=G_{i}({\mathbb{Q}}_{q}(\nu_{n}\omega_{n})/{\mathbb{Q}}_{q})\simeq G_{i}({\mathbb{Q}}_{q}(\nu_{n})/{\mathbb{Q}}_{q}).$
Moreover, again by Lemma 4.5 (iii), with $F={\mathbb{Q}}_{q}$ and with $\nu$, $\omega$ replaced by $\nu_{i}$, $\omega_{i}$ ($i=j,n$), and by (4.6)
@@ 2. $\bullet$
Then we proved points (i) and (ii).
To prove (iii), we recall that $\nu\omega\colon G_{{\mathbb{Q}}_{q}}\rightarrow{\mathbb{Z}}_{q}^{\times}$ is surjective by Lemma 4.5 (i). Scalar matrices with coefficients in ${\mathbb{Z}}_{p}^{\times}$ are central elements, invariant by the algebraic conjugation of ${\mathbb{Q}}_{q}/{\mathbb{Q}}_{p}$ and thus, by (4.5), are in the image of $\rho|_{G_{\mathbb{Q}_{q}}}$. ∎
@@
Let $p$ be a rational prime satisfying 3.3. Then the normal closure of $\mathrm{Gal}({\mathbb{Q}}_{q}(p^{n})/{\mathbb{Q}}_{q}(p^{n-1}))$ in $\mathrm{Gal}({\mathbb{Q}}(p^{n})/{\mathbb{Q}})$ is $\mathrm{Gal}({\mathbb{Q}}(p^{n})/{\mathbb{Q}}(p^{n-1}))$.
@@
{tikzpicture}
@@
$\widehat{G}(\mathcal{O}/p\mathcal{O})=\{\alpha\in\mathrm{GL}_{2}(\mathcal{O}/p\mathcal{O})\ |\ \det(\alpha)\in{\mathbb{F}}_{p}^{\times}\}.$
Assume first $n=1$. By Proposition 4.2, $\rho_{1}(D_{1})$ is a subgroup of $\mathrm{GL}_{2}({\mathbb{F}}_{p})$; by 4.7 (ii), its cardinality is $q-1>2(p-1)$. Moreover the determinant maps surjectively $\rho_{1}(D_{1})$ onto ${\mathbb{F}}_{p}^{\times}$, because its image is that of the cyclotomic character modulo $p$: $G_{\mathbb{Q}_{p}}\to\mathbb{F}_{p}^{\times}$. Thus $\rho_{1}(D_{1})$ satisfies the hypotheses of Proposition A.1.1, so that the normal closure of $\rho_{1}(D_{1})$ in $\widehat{G}(\mathcal{O}/p\mathcal{O})$ is $\widehat{G}(\mathcal{O}/p\mathcal{O})$ itself. Since by the hypothesis 3.3 (P2), $\rho_{1}$ is an isomorphism between $H_{1}$ and $\widehat{G}(\mathcal{O}/p\mathcal{O})$, the normal closure of $D_{1}$ in $H_{1}$ is $H_{1}$ itself.
Now we turn to the case $n\geq 2$. Let $\sigma\in H_{n}$. As in the proof of [15, Lemma 6.2], $\rho_{n}(\sigma)$ is the reduction mod $p^{n}$ of a matrix of the shape ${\rm I}+p^{n-1}A_{\sigma}$, where ${\rm I}$ denotes the $2$-by-$2$ identity matrix and where the “log” matrix $A_{\sigma}\in M_{2}(\mathcal{O})$ is well-defined mod $p\mathcal{O}$. Moreover, $\det(\rho_{n}(\sigma))\in({\mathbb{Z}}/p^{n}{\mathbb{Z}})^{\times}$ by Remark 3.2. Since $\det({\rm I}+p^{n-1}A_{\sigma})\equiv 1+{\rm trace}(A_{\sigma})p^{n-1}\!\!\!\mod p^{n}$, the matrix $A_{\sigma}$ has trace in ${\mathbb{F}}_{p}$. Let $\widehat{M}_{2}(\mathcal{O}/p\mathcal{O})$ be the subgroup of $\mathrm{M}_{2}(\mathcal{O}/p\mathcal{O})$ consisting of matrices having trace in ${\mathbb{F}}_{p}$. We denote by $\mathcal{L}_{n}$ the map $H_{n}\rightarrow\widehat{M}_{2}(\mathcal{O}/p\mathcal{O})$ which sends $\sigma$ to $A_{\sigma}$. Thus we have the following commutative diagram :
                 {tikzpicture}
Since $\rho_{n}$ and “log” are both injective, $\mathcal{L}_{n}$ is injective as well.
@@
where $\mathrm{M}_{2}(\mathcal{O}/p\mathcal{O})^{0}$ is the set of trace zero matrices. We also remark that $\mathcal{L}_{n}(D_{n})\subseteq\mathrm{M}_{2}({\mathbb{F}}_{p})$, by Proposition. As in the proof of [15, Lemma 6.2], for $\tau\in H_{n}$ and $\sigma\in\mathrm{Gal}({\mathbb{Q}}(p^{n})/{\mathbb{Q}})$ one has
@@
     {tikzpicture}
and “res” denotes the restriction. Let $\overline{D_{n}}$ be the normal closure of $D_{n}$ in $\mathrm{Gal}({\mathbb{Q}}(p^{n})/{\mathbb{Q}})$. Recall that $\overline{D_{n}}\subseteq H_{n}$. We want to show that $\mathcal{L}_{n}(\overline{D_{n}})=\widehat{M_{2}}(\mathcal{O}/p\mathcal{O})$. This will imply that $\overline{D_{n}}=H_{n}$ coincides with $H_{n}$ as we want (and that $\mathcal{L}_{n}$ is surjective). Note that, by property 3.3 (P2) $\rho_{1}$ is surjective. Since “res” is obviously surjective, $\tilde{\rho}_{n}$ is surjective as well. Thus (5.1) says in particular that $\mathcal{L}_{n}(\overline{D_{n}})$ is invariant by the adjoint action [^3] That is, the action by conjugation. of $\widehat{G}(\mathcal{O}/p\mathcal{O})$. Note that $\mathrm{M}_{2}(\mathcal{O}/p\mathcal{O})^{0}$ is also invariant by this action. Thus
$V=\mathcal{L}_{n}(\overline{D_{n}})\cap\mathrm{M}_{2}(\mathcal{O}/p\mathcal{O})^{0}$
@@
$V\cap\mathrm{M}_{2}({\mathbb{F}}_{p})^{0}=\mathcal{L}_{n}(\overline{D_{n}})\cap\mathrm{M}_{2}({\mathbb{F}}_{p})^{0}\neq\{0\}.$
As already remarked $\mathcal{L}_{n}$ of $D_{n}$ is contained in $\mathrm{M}_{2}({\mathbb{F}}_{p})$ by Proposition 4.2. Since $|D_{n}|=p^{2}$ by Lemma 4.7(i), its intersection with $\mathrm{M}_{2}({\mathbb{F}}_{p})^{0}$ is already not trivial. By Proposition A.2.5, $V=\mathrm{M}_{2}(\mathcal{O}/p\mathcal{O})^{0}$ and thus $\mathcal{L}_{n}(\overline{D_{n}})$ contains $\mathrm{M}_{2}(\mathcal{O}/p\mathcal{O})^{0}$.
The cyclotomic character $D_{n}\to(1+p^{n-1}\mathbb{Z}_{p})/(1+p^{n}\mathbb{Z}_{p})$ is surjective. Indeed by Proposition 4.1 and Mackey formula

### v3

@@ 2. Notations and preliminary results
    Let $j$ and $i$ be integers with $1\leq j\leq n$ and $q^{j-1}\leq(q-1,k-1)i\leq q^{j}-1$. The higher ramification groups are given by
    $G_{i}({\mathbb{Q}}_{q}(\epsilon_{2,n}^{k-1})/{\mathbb{Q}}_{q}))=\mathrm{Gal}({\mathbb{Q}}_{q}(\epsilon_{2,n}^{k-1})/{\mathbb{Q}}_{q}(\epsilon_{2,j}^{k-1})).$
Proof.
By [10, Théorème I.2.1], $\epsilon_{2}$ is up to isogeny the Tate module of a Lubin–Tate formal group $\mathcal{F}$ over ${\mathbb{Z}}_{q}$. Then ${\mathbb{Q}}_{q}(\epsilon_{2,n})={\mathbb{Q}}_{q}(\mathcal{F}(n))$, where $\mathcal{F}(n)$ is the group of $p^{n}$-division points of $\mathcal{F}$. Then in the case $k=2$ (i) and (ii) follow from [19, Theorem V.5.4] and [19, Proposition V.6.1].
Assume now $k>2$. The extension ${\mathbb{Q}}_{q}(\epsilon_{2,n}^{k-1})/{\mathbb{Q}}_{q}$ is the subextension of ${\mathbb{Q}}_{q}(\epsilon_{2,n})/{\mathbb{Q}}_{q}$ cut out by the $(k-1)$-th power of $\epsilon_{2,n}$. By assertion (i) in the special case $k=2$, and since $p\nmid k-1$ by Assumption 1.3 $(P3)$, we see that $\mathrm{Gal}({\mathbb{Q}}_{q}(\epsilon_{2,n})/{\mathbb{Q}}_{q}(\epsilon_{2,n}^{k-1}))$ is the subgroup of ${\mathbb{Z}}/(q-1){\mathbb{Z}}$ of index $d:=(q-1,k-1)$. Assertion (i) follows.
We now prove (ii). The group $\mathrm{Gal}({\mathbb{Q}}_{q}(\epsilon_{2,n})/{\mathbb{Q}}_{q}(\epsilon_{2,n}^{k-1}))$ has order $d$. By Herbrand Theorem [19, Ch. II, Theorem 10.7], $G_{r}({\mathbb{Q}}_{q}(\epsilon_{2,n})/{\mathbb{Q}}_{q})$ is mapped onto $G_{\eta(r)}({\mathbb{Q}}_{q}(\epsilon_{2,n}^{k-1})/{\mathbb{Q}}_{q})$ by the restriction map
$\mathrm{res}\colon\mathrm{Gal}({\mathbb{Q}}_{q}(\epsilon_{2,n})/{\mathbb{Q}}_{q})\longrightarrow\mathrm{Gal}({\mathbb{Q}}_{q}(\epsilon_{2,n}^{k-1})/{\mathbb{Q}}_{q}).$
Here (cf. [19, Ch. II, Proposition 10.6])
$\eta(r)=\frac{1}{d}\sum_{\sigma\in\mathrm{Gal}({\mathbb{Q}}_{q}(\epsilon_{2,n})/{\mathbb{Q}}_{q}(\epsilon_{2,n}^{k-1}))}\min\{i(\sigma),r+1\}-1$
with $i(\sigma)=v(\sigma(x)-x)$ for a generator $x$ of the ring of integers of ${\mathbb{Q}}_{q}(\epsilon_{2,n})$ over ${\mathbb{Q}}_{q}(\epsilon_{2,n})$. For $\sigma\in\mathrm{Gal}({\mathbb{Q}}_{q}(\epsilon_{2,n})/{\mathbb{Q}}_{q}(\epsilon_{2,n}^{k-1}))$ we have
$i(\sigma)=\begin{cases}\infty&\hbox{if }\sigma=id\\ 1&\hbox{ if }\sigma\not=id,\end{cases}$
because $G_{1}({\mathbb{Q}}_{q}(\epsilon_{2,n})/{\mathbb{Q}}_{q})$ is a $p$-group and $\mathrm{Gal}({\mathbb{Q}}_{q}(\epsilon_{2,n})/{\mathbb{Q}}_{q}(\epsilon_{2,n}^{k-1}))$ has order dividing $q-1$ and thus $G_{1}({\mathbb{Q}}_{q}(\epsilon_{2,n})/{\mathbb{Q}}_{q})\cap\mathrm{Gal}({\mathbb{Q}}_{q}(\epsilon_{2,n})/{\mathbb{Q}}_{q}(\epsilon_{2,n}^{k-1}))$ is trivial. Therefore
$\eta(r)=\frac{1}{d}(r+1+(d-1))-1=\frac{r}{d}.$
It follows that, for $j$ and $i$ integers such that $1\leq j\leq n$ and $q^{j-1}\leq d\,i\leq q^{j}-1$,
$G_{i}({\mathbb{Q}}_{q}(\epsilon_{2,n}^{k-1})/{\mathbb{Q}}_{q}))$ $=\mathrm{res}\left(G_{d\,i}({\mathbb{Q}}_{q}(\epsilon_{2,n})/{\mathbb{Q}}_{q})\right)$
$=\mathrm{res}\left({\mathbb{Q}}_{q}(\epsilon_{2,n})/{\mathbb{Q}}_{q}(\epsilon_{2,j})\right)\quad\hbox{(by the case $k=2$)}$
$=\mathrm{Gal}({\mathbb{Q}}_{q}(\epsilon_{2,n}^{k-1})/{\mathbb{Q}}_{q}(\epsilon_{2,j}^{k-1})).$
Now (ii) follows from (i). ∎
Proposition 4.7.
Let $n\geq 1$.
  1. (i)
    The extension ${\mathbb{Q}}_{q}(p^{n})/{\mathbb{Q}}_{q}$ is totally ramified and abelian of degree $\delta q^{n-1}$. Moreover
    $\mathrm{Gal}\left({\mathbb{Q}}_{q}(p^{n})/{\mathbb{Q}}_{q}\right)\simeq{\mathbb{Z}}/\delta{\mathbb{Z}}\times\left({\mathbb{Z}}/p^{n-1}{\mathbb{Z}}\right)^{2}$
  2. (ii)
    Let $j$ and $i$ be integers with $1\leq j\leq n$ and $q^{j-1}\leq(q-1,k-1)i\leq q^{j}-1$.The higher ramification groups are given by
@@
    In particular, the last non trivial higher ramification group has index
    $i_{n}:=\frac{q^{n-1}-1}{(q-1,k-1)}$
    and
    $G_{i_{n}}=\begin{cases}({\mathbb{Z}}/p{\mathbb{Z}})^{2}&\hbox{if }n\geq 2\\ {\mathbb{Z}}/\delta{\mathbb{Z}}&\hbox{if }n=1.\end{cases}$
@@
    If $M$ is an integer prime to $p$ and $v\mid p$, then the image of ${\rho_{v}}_{|G_{{\mathbb{Q}}_{q}}}$ contains the scalar matrix $\left(\begin{smallmatrix}M^{k-1}&0\\ 0&M^{k-1}\end{smallmatrix}\right)$.
@@
$\rho_{v|G_{{\mathbb{Q}}_{q}}}\simeq\left(\mathrm{ind}_{G_{{\mathbb{Q}}_{q}}}^{G_{{\mathbb{Q}}_{p}}}\epsilon_{2}^{k-1}\right)\otimes\mu_{\sqrt{-1}}.$
By Mackey’s restriction formula, we see that
(4.4) $\left(\mathrm{ind}_{G_{{\mathbb{Q}}_{q}}}^{G_{{\mathbb{Q}}_{p}}}\epsilon_{2}^{k-1}\right)_{|G_{{\mathbb{Q}}_{q}}}=\epsilon_{2}^{k-1}\oplus{\epsilon^{\prime}_{2}}^{k-1},$
where $\epsilon^{\prime}_{2}$ is the character obtained from $\epsilon_{2}$ by composing with the algebraic conjugation of ${\mathbb{Q}}_{q}/{\mathbb{Q}}_{p}$. Let for short $\nu=\epsilon_{2}^{k-1}$, $\omega=\mu_{\sqrt{-1}|G_{{\mathbb{Q}}_{q}}}$ and remark that $\omega$ takes values in {±$\{\pm 1\}$ and thus it is stable by the algebraic conjugation of ${\mathbb{Q}}_{q}/{\mathbb{Q}}_{p}$. Notice also that $-1\in{\mathrm{im}}(\nu)=\left(\mathbb{Z}_{q}^{\times}\right)^{k-1}$, because the weight $k$ is even. Then $\rho_{v|G_{{\mathbb{Q}}_{q}}}$ is conjugate to
(4.5) $G_{{\mathbb{Q}}_{q}}$ $\longrightarrow\mathrm{GL}_{2}({\mathbb{Z}}_{q})$
$\sigma$ $\to\left(\begin{smallmatrix}(\nu\omega)(\sigma)&0\\ 0&(\nu\omega)^{\prime}(\sigma)\end{smallmatrix}\right).$
@@
$\ker\rho(p^{n})_{|G_{{\mathbb{Q}}_{q}}}$ =$=\{\sigma\in G_{{\mathbb{Q}}_{q}}\ |\ (\nu\omega)(\sigma)\equiv(\nu\omega)^{\prime}(\sigma)\equiv 1\pmod{p^{n}}\}$
=$=\{\sigma\in G_{{\mathbb{Q}}_{q}}\ |\ (\nu\omega)(\sigma)\equiv 1\pmod{p^{n}}\}=\ker(\nu_{n}\omega_{n}),$
where we have denoted by $\nu_{n},\omega_{n}$, the composition of $\nu,\omega$ respectively with the reduction modulo $p^{n}$. Thus
@@
By its definition, $\nu_{n}$ is totally ramified; on the other hand $\omega$ is unramified with finite image contained in $\mathrm{im}(\nu)$. By (4.6), by Lemma 4.5 (iii) with $L={\mathbb{Q}}_{q}$ and with $\nu$, $\omega$ replaced by $\nu_{n}$, $\omega_{n}$ respectively, and by Lemma 4.6
$\mathrm{Gal}({\mathbb{Q}}_{q}(p^{n})/{\mathbb{Q}}_{q})=\mathrm{Gal}({\mathbb{Q}}_{q}(\nu_{n}\omega_{n})/{\mathbb{Q}}_{q})\simeq\mathrm{Gal}({\mathbb{Q}}_{q}(\nu_{n})/{\mathbb{Q}}_{q})\simeq{\mathbb{Z}}/\delta{\mathbb{Z}}\times\left({\mathbb{Z}}/p^{n-1}{\mathbb{Z}}\right)^{2}$
which proves (i), and
$G_{i}({\mathbb{Q}}_{q}(p^{n})/{\mathbb{Q}}_{q})=G_{i}({\mathbb{Q}}_{q}(\nu_{n}\omega_{n})/{\mathbb{Q}}_{q})\simeq G_{i}({\mathbb{Q}}_{q}(\nu_{n})/{\mathbb{Q}}_{q})=\mathrm{Gal}({\mathbb{Q}}_{q}(\nu_{n})/{\mathbb{Q}}_{q}(\nu_{j})).$
Moreover, again by Lemma 4.5 (iii), with $L={\mathbb{Q}}_{q}$ and with $\nu$, $\omega$ replaced by $\nu_{i}$, $\omega_{i}$ ($i=j,n$), and by (4.6)
@@
The last two displayed lines prove (ii).
To prove (iii), we recall that $\nu\omega\colon G_{{\mathbb{Q}}_{q}}\rightarrow({\mathbb{Z}}_{q}^{\times})^{k-1}$ is surjective by Lemma 4.5 (i). Scalar matrices with coefficients in $({\mathbb{Z}}_{p}^{\times})^{k-1}$ are central elements, invariant by the algebraic conjugation of ${\mathbb{Q}}_{q}/{\mathbb{Q}}_{p}$ and thus, by (4.5), are in the image of $\rho|_{G_{\mathbb{Q}_{q}}}$. ∎
@@
Let $p$ be a rational prime satisfying Assumptions 1.3. The normal closure of $\mathrm{Gal}({\mathbb{Q}}_{q}(p^{n})/{\mathbb{Q}}_{q}(p^{n-1}))$ in $\mathrm{Gal}({\mathbb{Q}}(p^{n})/{\mathbb{Q}})$ is $\mathrm{Gal}({\mathbb{Q}}(p^{n})/{\mathbb{Q}}(p^{n-1}))$.
@@
$H_{n}$$\mathrm{Gal}({\mathbb{Q}}_{q}(p^{n})/{\mathbb{Q}}_{q}(p^{n-1}))$$\Omega_{n}$$\mathrm{Gal}({\mathbb{Q}}(p^{n})/{\mathbb{Q}}(p^{n-1}))$$\mathrm{Gal}({\mathbb{Q}}(p^{n})/{\mathbb{Q}})$$\mathrm{GL}_{2}(\mathcal{O}/p^{n}\mathcal{O})$.$\bar{\rho}_{n}=\rho(p^{n})$
@@
Assume first $n=1$.
By Proposition 4.2, $\bar{\rho}_{1}(H_{1})$ is a subgroup of $\mathrm{GL}_{2}({\mathbb{F}}_{p})$; by Proposition 4.7 (ii), its cardinality is $\delta=\frac{q-1}{(q-1,k-1)}$. Let $s=\frac{p-1}{(p-1,k-1)}$. By Assumption 1.3 $(P3)$, $\frac{p+1}{2}\nmid k-1$. Thus $\frac{p+1}{(p+1,k-1)}>2$ and $\delta>2s$. The determinant maps surjectively $\bar{\rho}_{1}(H_{1})$ onto $({\mathbb{F}}_{p}^{\times})^{k-1}=({\mathbb{F}}_{p}^{\times})^{(k-1,p-1)}$, because its image is that of the $(k-1)$-th power $G_{\mathbb{Q}_{p}}\to(\mathbb{F}_{p}^{\times})^{k-1}$ of the cyclotomic character modulo $p$. Thus $\bar{\rho}_{1}(H_{1})$ satisfies the hypotheses of Proposition A.1.1, so that the normal closure of $\bar{\rho}_{1}(H_{1})$ in $\widehat{G}(\mathcal{O}/p\mathcal{O})$ is $\widehat{G}(\mathcal{O}/p\mathcal{O})$ itself. Since by the hypothesis 1.3 $(P2)$, $\bar{\rho}_{1}$ is an isomorphism between $G_{1}$ and $\widehat{G}(\mathcal{O}/p\mathcal{O})$, the normal closure of $H_{1}$ in $G_{1}$ is $G_{1}$ itself.
Now we turn to the case $n\geq 2$. Let $\sigma\in\Omega_{n}$. As in the proof of [15, Lemma 6.2], $\bar{\rho}_{n}(\sigma)$ is the reduction mod $p^{n}$ of a matrix of the shape ${\rm I}+p^{n-1}A_{\sigma}$, where ${\rm I}$ denotes the $2$-by-$2$ identity matrix and where the “log” matrix $A_{\sigma}\in M_{2}(\mathcal{O})$ is well-defined mod $p\mathcal{O}$. Moreover, $\det(\bar{\rho}_{n}(\sigma))\in({\mathbb{Z}}/p^{n}{\mathbb{Z}})^{\times}$ by Remark 2.2. Since $\det({\rm I}+p^{n-1}A_{\sigma})\equiv 1+{\rm trace}(A_{\sigma})p^{n-1}\!\!\!\mod p^{n}$, the matrix $A_{\sigma}$ has trace in ${\mathbb{F}}_{p}$. Let $\widehat{M}_{2}(\mathcal{O}/p\mathcal{O})$ be the subgroup of $\mathrm{M}_{2}(\mathcal{O}/p\mathcal{O})$ consisting of matrices having trace in ${\mathbb{F}}_{p}$. We denote by $\mathcal{L}_{n}$ the map $\Omega_{n}\rightarrow\widehat{M}_{2}(\mathcal{O}/p\mathcal{O})$ which sends $\sigma$ to $A_{\sigma}$. Thus we have the following commutative diagram :
                  $\Omega_{n}$${\rm Im}(\bar{\rho}_{n})$
$\subseteq$
${\rm I}+p^{n-1}M_{2}(\mathcal{O}/p^{n}\mathcal{O})$$\widehat{M}_{2}(\mathcal{O}/p\mathcal{O})$$\bar{\rho}_{n}$“log”$\mathcal{L}_{n}$
Since $\bar{\rho}_{n}$ and “log” are both injective, $\mathcal{L}_{n}$ is injective as well.
@@
where $\mathrm{M}_{2}(\mathcal{O}/p\mathcal{O})^{0}$ is the set of trace zero matrices. We also remark that $\mathcal{L}_{n}(H_{n})\subseteq\mathrm{M}_{2}({\mathbb{F}}_{p})$, by Proposition 4.2. As in the proof of [15, Lemma 6.2], for $\tau\in\Omega_{n}$ and $\sigma\in\mathrm{Gal}({\mathbb{Q}}(p^{n})/{\mathbb{Q}})$ one has
@@
      $\mathrm{Gal}({\mathbb{Q}}(p^{n})/{\mathbb{Q}})$$\mathrm{Gal}({\mathbb{Q}}(p)/{\mathbb{Q}})$$\widehat{G}(\mathcal{O}/p\mathcal{O})$$\{\alpha\in\mathrm{GL}_{2}(\mathcal{O}/p\mathcal{O})\ |\ \det(\alpha)\in(\mathbb{F}_{p}^{\times})^{k-1}\}$,“res”$\bar{\rho}_{1}$$\tilde{\rho}_{n}$
and “res” denotes the restriction. Let $\overline{H_{n}}$ be the normal closure of $H_{n}$ in $\mathrm{Gal}({\mathbb{Q}}(p^{n})/{\mathbb{Q}})$. Recall that $\overline{H_{n}}\subseteq\Omega_{n}$. We want to show that $\mathcal{L}_{n}(\overline{H_{n}})=\widehat{M_{2}}(\mathcal{O}/p\mathcal{O})$. This will imply that $\overline{H_{n}}=\Omega_{n}$ coincides with $\Omega_{n}$ as we want (and that $\mathcal{L}_{n}$ is surjective). Note that, by property 1.3 $(P2)$ $\bar{\rho}_{1}$ is surjective. Since “res” is obviously surjective, $\tilde{\rho}_{n}$ is surjective as well. Thus (5.1) says in particular that $\mathcal{L}_{n}(\overline{H_{n}})$ is invariant by the adjoint action [^5] That is, the action by conjugation. of $\widehat{G}(\mathcal{O}/p\mathcal{O})$. Note that $\mathrm{M}_{2}(\mathcal{O}/p\mathcal{O})^{0}$ is also invariant by this action. Thus
$V=\mathcal{L}_{n}(\overline{H_{n}})\cap\mathrm{M}_{2}(\mathcal{O}/p\mathcal{O})^{0}$
@@
$V\cap\mathrm{M}_{2}({\mathbb{F}}_{p})^{0}=\mathcal{L}_{n}(\overline{H_{n}})\cap\mathrm{M}_{2}({\mathbb{F}}_{p})^{0}\neq\{0\}.$
Indeed, $\mathcal{L}_{n}(H_{n})$ is contained in $\mathrm{M}_{2}({\mathbb{F}}_{p})$ by Proposition 4.2; since $|H_{n}|=p^{2}$ by Lemma 4.7(i), its intersection with $\mathrm{M}_{2}({\mathbb{F}}_{p})^{0}$ is already not trivial.
Then, by Proposition A.2.5, $V=\mathrm{M}_{2}(\mathcal{O}/p\mathcal{O})^{0}$ and thus $\mathcal{L}_{n}(\overline{H_{n}})$ contains $\mathrm{M}_{2}(\mathcal{O}/p\mathcal{O})^{0}$.
The cyclotomic character $H_{n}\to(1+p^{n-1}\mathbb{Z}_{p})/(1+p^{n}\mathbb{Z}_{p})$ is surjective. Indeed by Proposition 4.1 and Mackey formula